OpenStudy (anonymous):

can someone answer this quadric surface using completing the square ?thank you :) x^2 + 4y^2 - z^2 - 6x + 8y + 4z = 0

OpenStudy (amistre64):

get your like variables together and complete the squares .... which part is giving you pause?

OpenStudy (amistre64):

you can pull it apart into 3 quads if you like .... that might be less data to cull thru.

OpenStudy (anonymous):

i don't know how to solve this..i'm sorry

OpenStudy (amistre64):

do you know how to complete a square?

OpenStudy (anonymous):

it means that it will have a two factors in each variables..all in all it has a 6 factors?

OpenStudy (anonymous):

it is x^2 - 6x , 4y^2 + 8y and z^2 + 4z?

OpenStudy (amistre64):

lets go ahead and split this up into 3 different parts: x^2 + 4y^2 - z^2 - 6x + 8y + 4z = 0 x^2 -6x 4y^2 +8y -z^2 + 4z now that its organized better, you just work each one in its own right .... complete the square for each part

OpenStudy (amistre64):

the key is, do you know how to complete a square? :)

OpenStudy (anonymous):

ah.. its -z^2..

OpenStudy (anonymous):

i'm sorry i forgot how to complete a square..

OpenStudy (amistre64):

basically, you add and subtract a useful form of zero. square half the "x" term ..... for example: x^2 -6x ; the x term is -6, -6/2 = -3, (-3)^2 = 9. Lets add and subtract 9 x^2 -6x + 9 - 9; and condense the square (x-3)^2 - 9

OpenStudy (amistre64):

z will be trick only in that you have to account for the -

OpenStudy (anonymous):

x^2 -6x ; the x term is -6, how it become -6/2?

OpenStudy (amistre64):

we are trying to make a complete square, a perfect square. numbers of the form (a+b)^2 are perfect squares: (2+5)^2 = 49, (3+1)^2 = 16, ... we want to condense this thing back into some (a + b)^2 form, thats all. In order to do that we have to modify its form, but not its value. Since adding zero changes nothing ... its useful to us. BUT what form of zero? lets expand (a+b)^2 = a^2 +2ab + b^2 ; let a=x (x+b)^2 = x^2 +2bx + b^2 notice that the b^2 is missing above. and we know that the "x" term in the middle is equal to 2b. 2b = -6 b = -6/2 = -3 b^2 = 9

OpenStudy (amistre64):

so, adding and subtract 9; 9-9 = 0 right? will allow us to "complete" the square and alter its form ... not its value

OpenStudy (amistre64):

4y^2 +8y 2b = 8 b = 8/2 = 4 b^2 = 16, +16-16 = 0 4y^2 +8y + (16-16) therefore 4y^2 +8y = (2y +4)^2 - 16

OpenStudy (anonymous):

oh..i see. thanks i will solve this..

OpenStudy (anonymous):

btw do you know about contour plotting?

OpenStudy (amistre64):

i would need a little more information to see if i have contour plotting rolling around in my head someplace :)

OpenStudy (anonymous):

do you know what kind of contour plot is this if f(x,y) = \[\sqrt{x^2 + y^2}\] ?

OpenStudy (amistre64):

so the height, or z, is dependent on x and y inputs if we hold one input constant, we can generate a cross section plot and such, are some ideas ive got z = sqrt(x^2+y^2) z^2 = x^2 + y^2 is a circle, and so im gonna extrapolate that this is a sphere or hemisphere as a guess

OpenStudy (amistre64):

http://web.monroecc.edu/manila/webfiles/calcnsf/javacode/calcplot3d.htm might be useful :)

OpenStudy (amistre64):

well it use to work fine, i guess they changed some stuff about

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%28x%5E2%2By%5E2%29 hmmm, its conic not spherical. it comes to a point at the origin and makes bigger circles as z gets bigger and bigger

OpenStudy (anonymous):

our topic to this is about contour plotting..

OpenStudy (amistre64):



OpenStudy (amistre64):

im not up to date on what contour plotting is, this is mostly from an old, decripid memory :)

OpenStudy (amistre64):

for your private message z = k z = x^2 + y for what values of x^2 + y does k = the stated set of values

OpenStudy (amistre64):

x^2 + y = -2 when y=-x^2 -2